find the measure, in terms of x, of each side of a square if the area = x^-6x+9
Answer by lmeeks54(111)

Your problem is missing an exponent on the x^ term; but you probably meant, x^2. So I will go on from there.
. . .
The area for a square (or any rectangle) is the length of one side multiplied by an adjacent side. For a square, the sides are of equal length.
. . .
The easiest way to solve this problem in terms of x is to factor the quadratic equation:
. . .
x^2 - 6x + 9 = (x - 3)(x - 3) = (x - 3)^2
. . .
The length of any of the sides of the square, in terms of x, is = x - 3
. . .
Email me if you have any questions.